OpenStudy (anonymous):

mathdummy has a quick question what are the common factors that need to be canceled? 7/5 (5x+30/7) -6

OpenStudy (anonymous):

I know that the answer is x+6 .... but how/why

OpenStudy (anonymous):

\[\frac{ 7 }{ 5 } * (\frac{ 5x+30 }{ 7 }) - 6= 0\]

OpenStudy (anonymous):

Like that right?

OpenStudy (anonymous):

yes, sorta, this is part of a function , and I don't understand the cancel common factor part...

OpenStudy (anonymous):

May I see the whole function?

OpenStudy (anonymous):

sure... \[f(x)= \frac{ 7 }{ 5 } x -6 and g(x) = \frac{ 5x+30 }{ 7 }\]

OpenStudy (anonymous):

inserting g(x) in function of f(x)

OpenStudy (anonymous):

Hmm, okay then.

OpenStudy (anonymous):

sorry, forgot to put what we are going for in the long run...

OpenStudy (anonymous):

Ah, I got it..

OpenStudy (anonymous):

So, first you have this: 7/5 (5x+30/7) -6 to: 7/5 (5(x+6)/7) -6 (cancel 7s and 5s) So you have: (x+6)-6 becomes: x

OpenStudy (anonymous):

(f